Calculate the number of bit strings of length 5 or less.
1
Expert's answer
2020-11-10T13:04:11-0500

Let "a_1a_2...a_n" be a bit string of length "n". So, "a_i" can have two values for each "i\in\{1,2,...,n\}". Then by Multiplication Principle the number of bit string of length "n" is "2\cdot2\cdot ...\cdot 2=2^n".


Consequently,  the number of bit strings of length 5 or less is equal to


"2+2^2+2^3+2^4+2^5=2+4+8+16+32=62."


Answer: 62